Circular No. 5901

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian As-

trophysical Observatory, write: “GRO J0422+32, a possible black hole, is currently undergoing another unexpected and dramatic outburst. Its optical counterpart has brightened by 4 mag during the past two weeks at an average rate of 0.3 mag/day. CCD photometry obtained with the F. L. Whipple Observatory’s 1.2-m telescope gives the following R magnitudes: Nov. 21.52 UT, 19.27 ± 0.03; 24.21, 19.43 ± 0.09; 25.72, 19.16 ± 0.05; 26.19, 18.78 ± 0.07; Dec. 7.42, 15.45 ± 0.02; 8.20, 15.50 ± 0.02; 8.41, 15.38 ± 0.02. The data from the last two nights indicate that the object is now probably near its maximum for this outburst event. GRO J0422+32 was discovered in outburst on 1992 Aug. 5 (IAUC 5580); two subsequent, short-lived outbursts occurred on 1992 Dec. 20 (IAUC 5685) and 1993 Aug. 12 (IAUC 5842). It is unprecedented for an x-ray nova to have four outbursts within 16 months. Since the current outburst may also be short-lived, prompt observations at x-ray and γ-ray energies are strongly encouraged.”

RX J1940.2–1025 R. Staubert, S. Friedrich, and M. Bässgen, Astronomisches Institut,

Universität Tübingen; R. K. Sood and S. D. James, Australian Defence Force Academy; and W. Kollatschny, Universitätssternwarte Göttingen, report: “We have performed spectroscopic observations of the newly detected ROSAT source RX J1940.2–1025 near NGC 6814 (IAUC 5850, 5852) on Aug. 8, Sept. 6-7, and Oct. 2 and 26-30. The spectra show several Balmer and high-excitation He emission lines, confirming the nature as a cataclysmic variable; we further suggest that the object is an intermediate polar. The emission lines show significant wavelength shift with time. Radial velocities determined from Hα are consistent with a sinusoidal modulation, defining a spectroscopic period of 12120 ± 3 s (90-percent confidence for four free parameters) and a half-amplitude of ± 177 km/s. Heliocentric zero phase is at 1993 Sept. 6.899 UT. We note that periods determined from xray data taken with EXOSAT, Ginga, and ROSAT are generally longer (by 10-30 s), and so is the photometric period (12150 s, IAUC 5852). Even though formal uncertainties in the x-ray periods are sometimes of the same order as the differences from the spectroscopic period, it is conceivable that the system shows more than one period.”

1993 December 8 Daniel W. E. Green